# PRISE: Learning Temporal Action Abstractions as a Sequence Compression Problem

### v1

@@ Abstract
Temporal action abstractions, along with belief state representations, are a powerful knowledge sharing mechanism for sequential decision making. In this work, we propose a novel view that treats inducing temporal action abstractions as a sequence compression problem. To do so, we bring a subtle but critical component of LLM training pipelines – input tokenization via byte pair encoding (BPE) – to the seemingly distant task of learning skills of variable time span in continuous control domains. We introduce an approach called Primitive Sequence Encoding (PRISE) that combines continuous action quantization with BPE to learn powerful action abstractions. We empirically show that high-level skills discovered by PRISE from a multitask set of robotic manipulation demonstrations significantly boost the performance of both multitask imitation learning as well as few-shot imitation learning on unseen tasks.  [^1] Our code will be released at https://github.com/FrankZheng2022/PRISE.
@@ 1 Introduction
High-dimensional observations and decision-making over complex, continuous action spaces are hallmarks of typical scenarios in sequential decision-making, including robotics. A common way of dealing with these complexities is constructing abstractions – compact representations of belief states and actions that generalize across tasks and make learning to act in new scenarios robust and data-efficient. Inspired by techniques that have enabled ground-breaking progresses in computer vision (CV) and natural language processing (NLP) over the past decade, much of continuous control research has focused on learning abstractions from data rather than hand-crafting them. The lion’s share of these efforts study learning multi-task representations, which has analogies to representation learning in CV and NLP and has been tackled by adapting these fields’ models (see, e.g., DT (Chen et al., 2021) vs. GPT-2 (Radford et al., 2019)) and methods (see, e.g., R3M (Nair et al., 2022) vs. InfoNCE (den Oord et al., 2019)).
@@ 2 Preliminaries
We consider a set of tasks $T$, denoted as $\big{\{}(\mathcal{O},\mathcal{A},\mathcal{P},g_{\mathcal{T}})\big{\}}_{\mathcal{T}\in\mathscr{T}}$ that share the action space $\mathcal{A}$, the transition dynamics $\mathcal{P}$, and the observation space $\mathcal{O}$. The tasks differ only in their goals $g_{\mathcal{T}}$. Since we focus on learning visuomotor continuous control policies, we assume $\mathcal{A}$ is a finite-dimensional vector space and $\mathcal{O}$ is the image space.
@@
Byte Pair Encoding (BPE).
After quantizing the action space into codes, PRISE identifies common code (action) sequences via a technique from the realm of NLP, the method of Byte Pair Encoding (BPE) (Gage, 1994; Sennrich et al., 2016). In NLP, BPE has emerged as a pivotal technique for managing the vast and varied vocabulary encountered in textual data. Language models are expected to predict a text completion based on a text prefix. Doing so naively, by generating the text character-by-character is problematic, because the notion of a character differs significantly across languages. Instead, NLP models operate at the language-agnostic level of bytes. However, the high granularity of byte prediction has its own computational challenges. Therefore, language models models first find high-frequency byte sequences in the training data, assign tokens to these sequences, learn to predict the resulting tokens, which then get decoded into text. BPE is the algorithm that handles the common byte sequence construction problem.
BPE operates by merging the most frequent pairs of bytes, assigning tokens to these pairs, and iteratively merging new pairs of tokens or bytes. Thereby, it builds up a vocabulary of tokens that plays a crucial role in modern large language models OpenAI (2023); Brown et al. (2020); Radford et al. (2019). See Figure 2 for a demonstration. We apply BPE in the same way, but to action codes instead of bytes.
@@ 3.1 Pretraining
$\mathcal{L}_{\text{dyn}}[$ $\theta_{\mathcal{F},\mathcal{G},Q,P}](o_{t,T},a_{t})=-\textbf{cos\_sim}\big{(}Q(P(\mathcal{T}(z_{t},e_{t}))),\>\text{stopgrad}(P(z_{t+1}))\big{)}$ (1)
Here, $o_{t,T}=o_{t},o_{t-1},...,o_{t-T+1}$ is the observation history, $z_{t}=\mathcal{G}(o_{t},o_{t-1},...,o_{t-T+1})$ is the latent state, $e_{t}=\mathcal{F}(\text{stopgrad}(z_{t}),a_{t})$ is the quantized action code, $P$ is a projection MLP, and $Q$ is a prediction MLP. In our implementation, $P$ has 2 layers and $Q$ has 1.
Next, to guarantee the raw action can be decoded from the action code and latent state effectively, we train a latent state-dependent decoder, $\psi$, aiming to reconstruct the raw action given latent state $z_{t}$ and action code $e_{t}$. In general, we let $\psi$ be a stochastic mapping $\psi:\mathcal{Z}\times\mathcal{E}\rightarrow\Delta(\mathcal{A})$, where $\Delta(\mathcal{A})$ stands for the space of probability distributions over $\mathcal{A}$. We use a stochastic Gaussian Mixture Model (GMM) to parameterize the decoded action distribution of $\psi$ following Mandlekar et al. (2021). This choice has been found effective in dealing with the inherent multimodality and noise in such human teleoperation demonstrations.
Then the learning objective becomes to minimize the negative log likelihood of the GMM distribution:
$\mathcal{L}_{\text{act\_decode}}[\theta_{\mathcal{F},\mathcal{G},\psi}](o_{t,T},a_{t})=\mathcal{L}_{\text{act}}(\psi(z_{t},e_{t}),a_{t})=-\log\psi(a_{t}|z_{t},e_{t}).$ (2)
When we know the the pretraining data is collected by a fixed deterministic policy, we let $\psi:\mathcal{Z}\times\mathcal{E}\rightarrow\mathcal{A}$ be a deterministic function that minimizes the $L1$-loss:
$\mathcal{L}_{\text{act\_decode}}[\theta_{\mathcal{F},\mathcal{G},\psi}](o_{t,T},a_{t})=\mathcal{L}_{\text{act}}(\psi(z_{t},e_{t}),a_{t})=\|\psi(z_{t},e_{t})-a_{t}\|_{1}.$ (3)
@@
$\mathcal{L}[\theta_{\mathcal{F},\mathcal{G},Q,P,\psi}]=\mathcal{L}_{\text{dyn}}[\theta_{\mathcal{F},\mathcal{G},Q,P}]+\beta\mathcal{L}_{\text{act\_decode}}[\theta_{\mathcal{F},\mathcal{G},\psi}]$ (4)
@@ 3.3 Downstream few-shot adaptation to unseen tasks
$\mathcal{L}_{\textbf{FT\_DECODER}}[\theta_{\pi,\psi}]=\mathbb{E}_{\xi_{t}\sim\pi(\text{stopgrad}(z_{t}))}\Big{[}\mathcal{L}[\theta_{\psi}](\xi)\Big{]}$ (6)
@@
$\mathcal{L}[\theta_{\psi}](\xi_{t})=\sum_{i=0}^{\hat{K}}\mathcal{L}_{\text{action}}(\psi(\text{stopgrad}(z_{t+i}),\xi_{t}),a_{t+i})$ (7)
In this equation, $\hat{K}=\min(K,L_{\xi})$, where $K$ is a hyperparameter, the motivation behind which is explained at the end of Section 3.4.
@@
$\mathcal{L}[\theta_{\pi,\psi}]=\mathcal{L}_{\textbf{CE}}(\pi(\text{stopgrad}(z_{t})),\xi_{t})+\mathcal{L}_{\textbf{FT\_DECODER}}[\theta_{\pi,\psi}]\vspace{-0.8em}$ (8)
@@ 4 Related Work
Temporal Action Abstractions in Sequential Decision-Making: The concept is rooted in the literature of fully and partially observable MDPs (Puterman, 1994; Sutton & Barto, 2018). The complexity of policy learning is proportional to the problem horizon (Ross et al., 2011; Jiang et al., 2015; Rajaraman et al., 2020; Xie et al., 2021), leading to hierarchical approaches (Parr & Russell, 1998; Barto & Mahadevan, 2003; Le et al., 2018; Nachum et al., 2018; Kipf et al., 2019; Kujanpää et al., 2023) and the use of options or primitives (Sutton et al., 1999; Ajay et al., 2021).
@@
Temporal Abstractions vs. Skills in Robot Learning: Distinct from skill acquisition approaches like DMP (Pastor et al., 2009), Play-LMP (Lynch et al., 2019), and MCIL (Lynch & Sermanet, 2021), PRISE and methods like TAP (Jiang et al., 2023) focus on learning temporally extended action primitives.
Pretraining Data Assumptions: PRISE aligns more with Play-LMP (Lynch et al., 2019), RPL (Gupta et al., 2019), MCIL (Lynch & Sermanet, 2021), and TACO-RL (Rosete-Beas et al., 2022), requiring meaningful behavioral patterns in pretraining data.
@@ 5 Experiments
In this section, we empirically evaluate the effectiveness of the skill tokens of PRISE. We run PRISE to pretrain these tokens using large-scale, multitask offline datasets. Then we evaluate them on two offline IL scenarios: learning a multitask generalist policy and few-shot adaptation to unseen tasks. We show that PRISE is able to improve the performance of the learner policy, both compared to an PRISE version approaches that doesn’t use the skill tokens and compared to a very strong existing approach, ACT Zhao et al. (2023b).
@@ 5.1 Pretraining datasets and architecture
For LIBERO, we pretrain on the 90 short-horizon tasks (LIBERO-90) with offline dataset provided by the original paper. We test the learned skill tokens both in terms of the multitask learning performance on LIBERO-90 as well as 5-shot imitation learning (IL) performance on the first 8 unseen tasks from LIBERO-LONG. The pretraining dataset contains 50 demonstration trajectories for each task, collected by human teleoperation. We use the exact architecture of ResNet-T in Liu et al. (2023) for the observation embedding.
For MetaWorld, we focus on 5-shot IL since multitask IL on pretraining tasks is straightforward, with baseline algorithms, including PRISE, all achieving an average success rate of around 80%. We refer the readers to Figure 13 in Appendix D for the results. For the pretrain-test split, we hold out five hard tasks (hand-insert, box-close, disassemble, stick-pull, pick-place-wall) for few-shot evaluation and perform pretraining on the rest 45 tasks. We generate 100 expert trajectories for each pretraining task using the scripted policy provided in MetaWorld. Same as in as in Yarats et al. (2022), we use a shallow CNN encoder to encode the observation, and a transformer decoder module as the temporal backbone to encode temporal information into the latent embedding $z_{t}$.
For BPE on both domains, we set the codebook size $C$ to be 10 and vocabulary size to be 150, and we provide detailed ablation study to analyze their impacts later in this section. For more implementation details, we refer the readers to Appendix C.
Figure 6: Multitask policy learning performance on LIBERO-90. Error bar represents the standard deviation across 3 random seeds. Results of the first three methods are taken from (Liu et al., 2023)
@@ 5.2 Multitask generalist policy learning
Baselines. For multitask learning, our comparison includes PRISE and three network architectures for multitask behavior cloning (BC): ResNet-RNN, ResNet-T, and ViT-T, as implemented in (Liu et al., 2023). These architectures utilize either ResNet-18 (He et al., 2016) or ViT Dosovitskiy et al. (2021) to encode the pixel observations and then apply either a transformer or a LSTM module as temporal backbone to process a sequence of visual tokens. Recall that the main architecture of PRISE is the same as ResNet-T, except we add the vector quantization as well as other modules unique to PRISE. Additionally, we compare with ACT (Zhao et al., 2023a), an IL algorithm that learns a generative model for sequences of actions, termed “action chunks,” rather than single-step actions. These action chunks can be seen as a form of temporal action abstraction. During policy rollout, ACT combines overlapping action chunks predicted by its policy using a weighted average. Thus, the critical hyperparameter of ACT is the length of the action chunk, which we set at 8 after evaluating the best option from five choices: $\{3,5,8,10,15\}$.
@@ 5.3 Few-shot adaptation to unseen tasks
Figure 7: (Left): 5-shot IL performance averaged across 5 unseen MetaWorld tasks. (Right): 5-shot IL performance averaged across 8 unseen LIBERO tasks. Error bar stands for the standard deviation across 3 random seeds.
@@
Experimental Results (LIBERO). We present the average 5-shot IL performance across 8 tasks of LIBERO in LABEL:fig:libero_unseen. Different from MetaWorld, as shown in the figure, PRISE pretrained encoder does not improve the performance of the base IL algorithms for BC and ACT. This is consistent with what is reported in (Liu et al., 2023), where they observe that pretraining on LIBERO-90 with multitask BC even leads to negative transfer to downstream tasks. However, with pretrained skill tokens from PRISE, we could improve the average success rate by 9.2% compared with the best baseline algorithms. This further demonstrates the effectiveness of the proposed skill tokenization mechanisms by PRISE.

### v3

# PRISE: LLM-Style Sequence Compression for Learning Temporal Action Abstractions in Control

@@ Abstract
Temporal action abstractions, along with belief state representations, are a powerful knowledge sharing mechanism for sequential decision making. In this work, we propose a novel view that treats inducing temporal action abstractions as a sequence compression problem. To do so, we bring a subtle but critical component of LLM training pipelines – input tokenization via byte pair encoding (BPE) – to bear on the seemingly distant task of learning skills of variable time span in continuous control domains. We introduce an approach called Primitive Sequence Encoding (PRISE) that combines continuous action quantization with BPE to learn powerful action abstractions. We empirically show that high-level skills discovered by PRISE from a multitask set of robotic manipulation demonstrations significantly boost the performance of Behavior Cloning on downstream tasks.
Machine Learning, ICML
[…]
proof
@@ 1 Introduction
High-dimensional observations and decision-making over complex, continuous action spaces are hallmarks of typical scenarios in sequential decision-making, including robotics. A common way of dealing with these complexities is constructing abstractions – compact representations of belief states and actions that generalize across tasks and make learning to act in new scenarios robust and data-efficient. Inspired by techniques that have enabled ground-breaking progresses in computer vision (CV) and natural language processing (NLP) over the past decade, much of continuous control research has focused on learning abstractions from data rather than hand-crafting them. The lion’s share of these efforts study learning multi-task representations, which has analogies to representation learning in CV and NLP and has been tackled by adapting these fields’ models (see, e.g., DT (Chen et al., 2021) vs. GPT-2 (Radford et al., 2019)) and methods (see, e.g., R3M (Nair et al., 2022) vs. InfoNCE (den Oord et al., 2019; Zheng et al., 2023a)).
@@ 2 Preliminaries
We consider a set of tasks $T$, denoted as $\big{\{}(\mathcal{O},\mathcal{A},\mathcal{P},g_{\mathcal{T}})\big{\}}_{\mathcal{T}\in\mathscr{T}}$, that share the action space $\mathcal{A}$, the transition dynamics $\mathcal{P}$, and the observation space $\mathcal{O}$. The tasks differ only in their goals $g_{\mathcal{T}}$. Since we focus on learning visuomotor continuous control policies, we assume $\mathcal{A}$ is a finite-dimensional vector space and $\mathcal{O}$ is the image space.
@@
Byte Pair Encoding (BPE). After quantizing the action space into codes, PRISE identifies common code (action) sequences via a technique from the realm of NLP, the method of Byte Pair Encoding (BPE) (Gage, 1994; Sennrich et al., 2016). In NLP, BPE has emerged as a pivotal technique for managing the vast and varied vocabulary encountered in textual data. Language models are expected to predict a text completion based on a text prefix. Doing so naively, by generating the text character-by-character is problematic, because the notion of a character differs significantly across languages. Instead, NLP models operate at the language-agnostic level of bytes. However, the high granularity of byte prediction has its own computational challenges. Therefore, language models models first find high-frequency byte sequences in the training data, assign tokens to these sequences, learn to predict the resulting tokens, which then get decoded into text. BPE is the algorithm that handles the common byte sequence construction problem.
BPE operates by merging the most frequent pairs of bytes, assigning tokens to these pairs, and iteratively merging new pairs of tokens or bytes. Thereby, it builds up a vocabulary of tokens that plays a crucial role in modern large language models (OpenAI, 2023; Brown et al., 2020; Radford et al., 2019). See Figure 2 for a demonstration. We apply BPE in the same way, but to action codes instead of bytes.
@@ 3.1 Pretraining
$\mathcal{L}_{\text{dyn}}[$ $\theta_{\mathcal{F},\mathcal{G},Q,P}](o_{t,T},a_{t})=$ (1)
[…]
Here, $o_{t,T}=o_{t},o_{t-1},...,o_{t-T+1}$ is the observation history, $z_{t}=\mathcal{G}(o_{t},o_{t-1},...,o_{t-T+1})$ is the latent state, $e_{t}=\mathcal{F}(\text{stopgrad}(z_{t}),a_{t})$ is the quantized action code, $P$ is a projection MLP, and $Q$ is a prediction MLP. In our implementation, $P$ has two layers and $Q$ has one.
Next, to guarantee the raw action can be decoded from the action code and latent state effectively, we train a latent state-dependent decoder, $\psi$, aiming to reconstruct the raw action given latent state $z_{t}$ and action code $e_{t}$. In general, we let $\psi$ be a stochastic mapping $\psi:\mathcal{Z}\times\mathcal{E}\rightarrow\Delta(\mathcal{A})$, where $\Delta(\mathcal{A})$ stands for the space of probability distributions over $\mathcal{A}$. We use a stochastic Gaussian Mixture Model (GMM) to parameterize the decoded action distribution of $\psi$ following Mandlekar et al. (2021). This choice has been found effective in dealing with the inherent multimodality and noise in such human teleoperation demonstrations. Then the learning objective becomes to minimize the negative log likelihood of the GMM distribution: $\mathcal{L}_{\text{act\_decode}}[\theta_{\mathcal{F},\mathcal{G},\psi}](o_{t,T},a_{t})=\mathcal{L}_{\text{act}}(\psi(z_{t},e_{t}),a_{t})=-\log\psi(a_{t}|z_{t},e_{t})$. When we know the the pretraining data is collected by a fixed deterministic policy, we let $\psi:\mathcal{Z}\times\mathcal{E}\rightarrow\mathcal{A}$ be a deterministic function that minimizes the $L1$-loss: $\mathcal{L}_{\text{act\_decode}}[\theta_{\mathcal{F},\mathcal{G},\psi}](o_{t,T},a_{t})=\mathcal{L}_{\text{act}}(\psi(z_{t},e_{t}),a_{t})=\|\psi(z_{t},e_{t})-a_{t}\|_{1}.$
@@
$\mathcal{L}[\theta_{\mathcal{F},\mathcal{G},Q,P,\psi}]=\mathcal{L}_{\text{dyn}}[\theta_{\mathcal{F},\mathcal{G},Q,P}]+\beta\mathcal{L}_{\text{act\_decode}}[\theta_{\mathcal{F},\mathcal{G},\psi}]$
@@ 3.3 Downstream few-shot adaptation to unseen tasks
$\mathcal{L}_{\textbf{FT\_DECODER}}[\theta_{\pi,\psi}]=\mathbb{E}_{\xi_{t}\sim\pi(\text{stopgrad}(z_{t}))}\Big{[}\mathcal{L}[\theta_{\psi}](\xi_{t})\Big{]}$
@@
$\mathcal{L}[\theta_{\psi}](\xi_{t})=\sum_{i=0}^{\hat{K}}\mathcal{L}_{\text{action}}(\psi(\text{stopgrad}(z_{t+i}),\xi_{t}[i]),a_{t+i})$
In this equation, $\hat{K}=\min(K,L_{\xi})$, where $K$ is a hyperparameter, the motivation behind which is explained at the end of Section 3.4. $\xi_{t}[i]$ represents the $i$-th code of the skill token $\xi_{t}$.
@@
$\mathcal{L}[\theta_{\pi,\psi}]=\mathcal{L}_{\textbf{CE}}(\pi(\text{stopgrad}(z_{t})),\xi_{t})+\mathcal{L}_{\textbf{FT\_DECODER}}[\theta_{\pi,\psi}]$
@@ 4 Related Work
Temporal Action Abstractions in Sequential Decision-Making: Temporal action abstraction and options or skills by compression is not newly introduced in this paper. In fact, the concept of temporal action abstraction is rooted in the literature of both fully and partially observable MDPs (Puterman, 1994; Thrun & Schwartz, 1994; Sutton & Barto, 2018). The complexity of policy learning is proportional to the problem horizon (Ross et al., 2011; Jiang et al., 2015; Rajaraman et al., 2020; Xie et al., 2021), leading to hierarchical approaches (Parr & Russell, 1998; Barto & Mahadevan, 2003; Le et al., 2018; Nachum et al., 2018; Kipf et al., 2019; Kujanpää et al., 2023; Jiang et al., 2022) and the use of options or primitives (Sutton et al., 1999; Ajay et al., 2021; Cui et al., 2023).
@@
Temporal Abstractions vs. Skills in Robot Learning: Distinct from skill acquisition approaches such as LOVE (Jiang et al., 2022), BeT (Shafiullah et al., 2022), C-BeT (Cui et al., 2023), DMP (Pastor et al., 2009), Play-LMP (Lynch et al., 2019), and MCIL (Lynch & Sermanet, 2021), PRISE and methods like TAP (Jiang et al., 2023) focus on learning temporally extended action primitives.
Pretraining Data Assumptions: PRISE aligns more with C-BeT (Cui et al., 2023), Play-LMP (Lynch et al., 2019), RPL (Gupta et al., 2019), MCIL (Lynch & Sermanet, 2021), and TACO-RL (Rosete-Beas et al., 2022), requiring meaningful behavioral patterns in pretraining data.
@@ 5 Experiments
In this section, we empirically evaluate the effectiveness of the skill tokens of PRISE. We run PRISE to pretrain these tokens using large-scale, multitask offline datasets. Then we evaluate them on two offline IL scenarios: learning a multitask generalist policy and few-shot adaptation to unseen tasks. We show that PRISE is able to improve the performance of the learner policy, both compared to an PRISE version approaches that doesn’t use the skill tokens and compared to a very strong existing approach, ACT (Zhao et al., 2023b).
@@ 5.1 Pretraining datasets and architecture
For LIBERO, we pretrain on the 90 short-horizon tasks (LIBERO-90) with offline dataset provided by the original paper. We test the learned skill tokens both in terms of the multitask learning performance on LIBERO-90 as well as 5-shot imitation learning (IL) performance on the first 8 unseen tasks from LIBERO-LONG. The pretraining dataset contains 50 demonstration trajectories for each task, collected by human teleoperation. We use the exact architecture of ResNet-T in (Liu et al., 2023) for the observation embedding.
For MetaWorld, we focus on 5-shot IL since multitask IL on pretraining tasks is straightforward, with baseline algorithms, including PRISE, all achieving an average success rate of around 80%. We refer the readers to Figure 15 in Appendix D for the results. For the pretrain-test split, we hold out five hard tasks (hand-insert, box-close, disassemble, stick-pull, pick-place-wall) for few-shot evaluation and perform pretraining on the rest 45 tasks. We generate 100 expert trajectories for each pretraining task using the scripted policy provided in MetaWorld. Same as in as in (Yarats et al., 2022; Xu et al., 2024), we use a shallow CNN encoder to encode the observation, and a transformer decoder module as the temporal backbone to encode temporal information into the latent embedding $z_{t}$.
For BPE on both domains, we set the codebook size $C$ to be 10 and vocabulary size to be 200, and we provide detailed ablation study to analyze their impacts later in this section. For more implementation details, we refer the readers to Appendix C.
Figure 6: Multitask policy learning performance on LIBERO-90. Error bar represents the standard deviation across 4 random seeds. Results of the first three methods are taken directly from (Liu et al., 2023)
@@ 5.2 Multitask generalist policy learning
Baselines. For multitask learning, our comparison includes PRISE and three network architectures for multitask behavior cloning (BC): ResNet-RNN, ResNet-T, and ViT-T, as implemented in (Liu et al., 2023). These architectures utilize either ResNet-18 (He et al., 2016) or ViT (Dosovitskiy et al., 2021) to encode the pixel observations and then apply either a transformer or a LSTM module as temporal backbone to process a sequence of visual tokens. Recall that the main architecture of PRISE is the same as ResNet-T, except we add the vector quantization as well as other modules unique to PRISE. Additionally, we compare with ACT (Zhao et al., 2023a), an IL algorithm that learns a generative model for sequences of actions, termed “action chunks,” rather than single-step actions. These action chunks can be seen as a form of temporal action abstraction. During policy rollout, ACT combines overlapping action chunks predicted by its policy using a weighted average. Thus, the critical hyperparameter of ACT is the length of the action chunk, which we set at 8 after evaluating the best option from five choices: $\{3,5,8,10,15\}$.
@@ 5.3 Few-shot adaptation to unseen tasks
Figure 7: (Metaworld): 5-shot IL performance averaged across 5 unseen tasks. Error bar represents the standard deviation across 4 random seeds.
@@
Experimental Results (LIBERO). We present the average 5-shot IL performance across 8 tasks of LIBERO in Figure 8. Different from MetaWorld, as shown in the figure, PRISE pretrained encoder does not improve the performance of the base IL algorithms for BC and ACT. This is consistent with what is reported in (Liu et al., 2023), where they observe that pretraining on LIBERO-90 with multitask BC even leads to negative transfer to downstream tasks. However, with pretrained skill tokens from PRISE, we could improve the average success rate by 9.2% compared with the best baseline algorithms. This further demonstrates the effectiveness of the proposed skill tokenization mechanisms by PRISE.
Figure 8: (LIBERO): 5-shot IL performance averaged across 8 unseen tasks. Error bar stands for the standard deviation across 4 random seeds.